OpenStudy (anonymous):

I just need to know how to factor out the 9's and why when you factor out the 9's it becomes 9(3ln(2)+ln(3))

OpenStudy (anonymous):

9ln(2)+9ln(12)

OpenStudy (amistre64):

n(a+b) = na + nb

OpenStudy (amistre64):

its just multiplication; and to factor it out you simply divide it off

OpenStudy (amistre64):

the: 3ln(2)+ln(3) : part tho, i cant make sense of that as of yet

OpenStudy (amistre64):

ln(2) + ln(12) = ln(2*12) ln(24) = ln(8*3) ln(8*3) = ln(2^3) + ln(3) ln(8*3) = 3ln(2) + ln(3)

OpenStudy (phi):

For a short video on the distributive property, see http://www.khanacademy.org/video/the-distributive-property?playlist=Developmental+Math

OpenStudy (amistre64):

they just seem to be manipulating the ln properties to "simplify" things

OpenStudy (anonymous):

yea, if you check out the previous post, two people replyed to this question already